The 2nd term of a geometric series is 80 and the sixth term is 16/125. Find the common ratio and the first term of the series. Find also the sum to infinity of the series.
Answer by Cromlix(4381)

Consider the series as starting with 80
until you find the common ratio.
a1 = 80, a5 = 16/125 and n = 5
An = A2* r^5-1
16/125 = 80* r^4
Divide both sides by 80
16/10000 = r^4
Take the root 4 of 16/10000
= 2/10
= 1/5
So the common ratio is 1/5
Multiply 80 by 5
First term = 400.
....................
Sum to Infinity = a/(1 - r)
= 400/ (1 - 1/5)
=400/ 4/5
= 500
Hope this helps.
:-)